OpenStudy (anonymous):

how many moles of oxygen gas are needed to react with 4 moles c2h6?

OpenStudy (anonymous):

First you have to know the reactants and balance that equation. \[2C _{2}H _{6}+7O _{2} \rightarrow 4CO _{2}+6H _{2}O\] So the mole to mole formula is simple. \[Known Mole * \frac{ Unknown Mole Coefficient }{ Known Mole Coefficient }=Unknown Mole\] Ratio of unknown to known is \[\frac{ 7 }{ 2 }\] Just plug everything in and solve it: \[4 * \frac{ 7 }{ 2 } = Moles of O_{2}\] \[14 = Moles of O_{2}\]